OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS. Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS. @Mathematics

OpenStudy (eyust707):

haha theres a millions things you could do

OpenStudy (anonymous):

batman symbol + pi= santa

OpenStudy (anonymous):

that has no solutions

OpenStudy (eyust707):

yea it does: bat symbol = santa - pi

OpenStudy (eyust707):

=P

OpenStudy (anonymous):

an actual answer ?

OpenStudy (eyust707):

x^2 + 6 = -5

OpenStudy (eyust707):

x^4 + 999999 - 5 + 5 + sqrt6 = -33333333333

OpenStudy (eyust707):

ln -420 + 5 (7x) = x

OpenStudy (eyust707):

5x=7x

OpenStudy (eyust707):

x=2x -3x+9x

OpenStudy (anonymous):

why are there no solutions?

OpenStudy (eyust707):

hehe ok lets look at an easy one

OpenStudy (eyust707):

x = 2x + 3x

OpenStudy (eyust707):

simply that

OpenStudy (eyust707):

what do you get?

OpenStudy (anonymous):

magic

OpenStudy (eyust707):

maybe with a little..